# RestoreVAR: Visual Autoregressive Generation for All-in-One Image Restoration

### v1

@@ Abstract
The use of latent diffusion models (LDMs) such as Stable Diffusion has significantly improved the perceptual quality of All-in-One image Restoration (AiOR) methods, while also enhancing their generalization capabilities. However, these LDM-based frameworks suffer from slow inference due to their iterative denoising process, rendering them impractical for time-sensitive applications. To address this, we propose RestoreVAR, a novel generative approach for AiOR that significantly outperforms LDM-based models in restoration performance while achieving over $\mathbf{10\times}$ faster inference. RestoreVAR leverages visual autoregressive modeling (VAR), a recently introduced approach which performs scale-space autoregression for image generation. VAR achieves comparable performance to that of state-of-the-art diffusion transformers with drastically reduced computational costs. To optimally exploit these advantages of VAR for AiOR, we propose architectural modifications and improvements, including intricately designed cross-attention mechanisms and a latent-space refinement module, tailored for the AiOR task. Extensive experiments show that RestoreVAR achieves state-of-the-art performance among generative AiOR methods, while also exhibiting strong generalization capabilities. Project page: https://sudraj2002.github.io/restorevarpage/
@@ 1 Introduction
Image restoration is a complex inverse problem that aims to recover clean images from various forms of degradations, such as haze, rain, snow, blur, and low-light conditions. Recently, the paradigm of All-in-One image Restoration (AiOR) has emerged, where a single deep network is trained to handle multiple degradation types. Existing AiOR methods can be broadly categorized into non-generative and generative approaches. Non-generative models such as AirNet airnet , PromptIR promptir , InstructIR instructir , AWRaCLe awracle , and AdaIR adair , deterministically map degraded images to their clean counterparts. While these methods offer fast inference and reliable pixel-level restoration performance, they often fail to generalize to diverse degradations encountered in real-world scenarios. To overcome this challenge, recent works have adopted generative models that aim to capture the distribution of clean images and produce more perceptually realistic outputs. Early works gan1rain ; gan2rain ; gan3blur based on GANs gan attempted this through adversarial learning, but suffered from mode collapse and unstable training. To improve fidelity and training stability, methods such as DiffUIR diffuir and DA-CLIP daclip employed pixel-space diffusion models ddpm . However, their high computational cost makes large-scale pretraining infeasible, limiting their ability to learn strong generative priors. In contrast, recent methods such as Diff-Plugindiffplugin , AutoDIRautodir , and PixWizard pixwizard leverage latent diffusion models (LDMs), such as Stable Diffusion stablediff . By operating in a lower-dimensional latent space, LDMs significantly reduce computational costs, enabling large-scale pretraining which equips them with strong generative priors of natural images. These priors allow LDM-based AiOR methods to deliver perceptually realistic restoration and improved generalization to real-world degradations.
@@
Despite their advantages, LDM-based AiOR methods have some shortcomings. (1) LDMs require multiple denoising steps during inference, resulting in significantly longer runtimes compared to non-generative models. Their slow inference speeds pose challenges for applications that demand real-time processing, such as video surveillance or autonomous navigation. (2) LDMs rely on variational autoencoders (VAEs) vae which are primarily trained for generative diversity, rather than accurate pixel-level reconstruction. Consequently, the restored images obtained from LDM-based AiOR methods exhibit loss of fine structural details, hindering their performance.
@@
An alternative class of generative models are autoregressive models, which have driven rapid advances in natural language processing through large language models (LLMs) such as GPT-3 radford2019language and LLaMA touvron2023llama . These models generate outputs by predicting the next token in a sequence, conditioned on previously generated tokens. Recently, Visual AutoRegressive (VAR) Modeling var introduced a scale-space autoregressive framework for image generation, performing next-scale prediction in the latent space of a multi-scale vector-quantized VAE (VQVAE). VAR achieves performance comparable to those of state-of-the-art diffusion models such as DiT-XL/2 dit , while operating 45×45$\times$45 × faster. Despite its success in generative tasks, the application of VAR to low-level vision tasks such as image restoration remains largely unexplored. To the best of our knowledge, only two prior works—VarSR varsr and Varformer varformer —have attempted to use VAR for image restoration. VarSR focused exclusively on the super-resolution task, while Varformer utilized intermediate VAR features to guide a separate non-generative network for AiOR. In contrast, our approach is generative and fully exploits the strong priors of the pretrained VAR model by training it directly for the AiOR task.
@@
Figure 1: RestoreVAR, our proposed VAR-based var scale-space generative AiOR model (a), significantly outperforms LDM-based methods as shown in (b). RestoreVAR also offers drastic reductions in computational complexity as shown in (c).
To this end, we introduce RestoreVAR, a novel generative approach for AiOR that addresses some of the key limitations of LDM-based approaches. Firstly, RestoreVAR adopts the autoregressive structure of VAR, achieving state-of-the-art generative AiOR performance with over $\mathbf{10\times}$ faster inference than LDM-based methods (see Fig. 1). Secondly, RestoreVAR employs cross-attention mechanisms conditioned on the degraded image latents, enabling the model to maintain spatial consistency and minimize hallucinations. Thirdly, to mitigate the loss of fine details by the vector quantization and VAE decoding processes, we propose a lightweight (only $\sim\mathbf{3}$overhead) non-generative latent refinement transformer which predicts de-quantized latents from the outputs of VAR. Additionally, we fine-tune the VAE decoder to operate on these continuous latents, further enhancing reconstruction quality. Finally, through extensive experiments, we demonstrate that RestoreVAR achieves state-of-the-art performance among generative restoration models, while also exhibiting strong generalization to real-world degradations. To summarize, our key contributions are:
@@
    We propose RestoreVAR, a novel generative AiOR framework that achieves superior performance and a $\mathbf{10\times}$ faster inference than LDM-based methods. To the best of our knowledge, RestoreVAR is the first method to directly utilize the VAR model for the AiOR task.
@@ 2 Related Works
We now discuss prior works on image restoration and autoregressive models in computer vision.
@@ 2.1 Image restoration
Early restoration models primarily addressed specific degradations  early1 ; haze1 ; early3 ; rain1 ; rainnew1 ; rainnew2 ; snow1 ; snow2 ; gopro ; hide . Later methods such as Restormer restormer , MPRNet mprnet and SwinIR swinir introduced baseline architectures for any single restoration task. However, they are restricted to handle one degradation at a time, making them ineffective for multiple degradations. All-in-One image Restoration (AiOR) methods aim to tackle multiple corruptions with a single model. Early approaches include non-generative models such as All-in-one nas , which used neural architecture search, and Transweather transw for adverse weather restoration. PromptIR promptir used learnable prompts while AWRaCLe awracle utilized visual in-context learning to extract degradation characteristics. Other approaches such as InstructIR instructir adopted textual guidance, and DCPT dcpt proposed a novel pre-training strategy for AiOR. Recent AiOR methods have adopted diffusion models. Pixel-space diffusion models (PSDMs) such as DA-CLIP daclip and DiffUIR diffuir demonstrated improved AiOR performance. However, PSDMs lack the robust generative priors of latent diffusion models (LDMs) such as Stable Diffusion stablediff . Thus, recent methods have utilized the strong priors of LDMs for AiOR. Diff-Plugin diffplugin adopts task plugins to guide an LDM for AiOR. AutoDIR autodir developed an automatic approach for degradation detection and restoration using an LDM. PixWizard pixwizard is a multi-task SD-XL sdxl based model capable of performing AiOR among other tasks. However, LDM-based approaches are slow at inference time—a limitation we aim to overcome using visual autoregressive modeling (VAR) var .
@@ 2.2 Autoregressive models in vision
Autoregressive (AR) models generate sequences by predicting the next element conditioned on preceding ones, modeling the joint probability of the sequence as a product of conditional probabilities. This paradigm has revolutionized natural language processing (NLP) and forms the foundation of large language models (LLMs) such as GPT-3 radford2019language and LLaMA touvron2023llama . Recent works van2016pixel ; var have extended AR models to vision and can be categorized as pixel-space AR van2016pixel ; van2016conditional ; reed2017parallel ; chen2019pixelsnail , token-based AR oord2018neural ; yu2023video ; nash2021generating ; ramesh2021zeroshot and scale-space AR var ; ren2024mvar ; ren2024flowar ; guo2025fastvar . Pixel-space AR predicts raw pixels one by one in raster order, as in PixelRNN van2016pixel and PixelCNN++ salimans2017pixelcnn++ , but is prohibitively slow at high resolutions. Token-based AR compresses images into discrete latent codes via vector quantization (e.g., VQ-VAE oord2018neural , VQ-VAE-2 razavi2019generating , VQGAN esser2021taming ) and then models code sequences with transformers (e.g., Image Transformer parmar2018image , ImageGPT chen2020image ). This trades-off codebook size and transformer capacity against tractability for high-resolution generation. Scale-space AR, as introduced in VAR var , generates latents hierarchically from coarse to fine scales and matches the quality of Diffusion Transformers dit at a fraction of the inference costvar . HART hart scales VAR to $1024\times 1024$ synthesis by using a lightweight MLP-based diffusion refiner to convert discrete VAR latents into continuous representations. Despite VAR’s success in generative tasks, it remains underexplored for image restoration. Only two prior works-VarSR varsr and Varformer varformer -have exploited VAR’s generative priors for image restoration. VarSR addressed super-resolution, while Varformer used VAR’s features to guide a non-generative AiOR model. In contrast, RestoreVAR is a generative model which directly trains VAR for multiple image restoration tasks.
@@ 3 Proposed Method
In this section, we begin by explaining the working principles behind VAR for image generation. We then detail RestoreVAR, our proposed VAR-based approach for AiOR.
@@ 3.1 Preliminaries: Visual Autoregressive Modelling
Visual Autoregressive Modelling, or VAR, is a novel autoregressive class-conditioned image generation method which uses a GPT-2 gpt2 style decoder-only transformer architecture for next-scale prediction. The VAR transformer operates in the latent space of a multi-scale VQVAE which uses $K$ scales. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, the VQVAE encoder outputs a latent representation $f_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. Hereafter, we will refer to $f_{\text{cont}}$ as the continuous latent, and the latent obtained after quantization as discrete latent. Instead of directly quantizing $f_{\text{cont}}$, a multi-scale residual quantization using a shared codebook across $K$ spatial scales is performed. First, the residual and accumulated quantized (or discrete) reconstruction of $f_{\text{cont}}$ are initialized as $f_{\text{res}}^{(0)}:=f_{\text{cont}}$ and $f_{\text{quant}}^{(0)}:=0$, respectively. At each scale $k=1,\ldots,K$, an index map $r_{k}\in\mathbb{Z}^{H_{k}\times W_{k}}$ is obtained by quantizing the downsampled residual feature:
$r_{k}:=\text{quantize}\left(\text{downsample}\left(f_{\text{res}}^{(k-1)}\right)\right).$
@@
$h_{k}:=\phi_{k}\left(\text{upsample}\left(e(r_{k})\right)\right),\in\mathbb{R}^{H_{K}\times W_{K}\times C}.$
@@
$f_{\text{quant}}^{(k)}:=f_{\text{quant}}^{(k-1)}+h_{k},\quad f_{\text{res}}^{(k)}:=f_{\text{cont}}-f_{\text{quant}}^{(k)}.$
@@
$p(r_{1},r_{2},\ldots,r_{K})=\prod_{k=1}^{K}p(r_{k}\mid r_{1},r_{2},\ldots,r_{k-1}).$
@@ 3.2 RestoreVAR
We now describe RestoreVAR, our proposed approach that adapts the generative capabilities of VAR for AiOR, while leveraging its substantial inference speed advantage over LDMs. Given a degraded image $I_{\text{deg}}\in\mathbb{R}^{H\times W\times 3}$, the goal is to predict a clean output $I_{\text{clean}}$, close to the ground-truth $I_{\text{gt}}$. Adapting VAR to AiOR is non-trivial due to the need for high-quality pixel-level reconstruction, which is compromised by two factors: (1) VAR’s strong generative priors can cause hallucinations in the restored images without proper conditioning. (2) Vector quantization and VAE decoding introduce artifacts that hinder pixel-level restoration. RestoreVAR addresses these challenges through architectural enhancements, including cross-attention to incorporate semantic guidance from the degraded image, and a novel non-generative transformer that refines discrete latents into their continuous form to preserve fine details in the restored image. We describe these components below.

### v2

@@ Abstract
The use of latent diffusion models (LDMs) such as Stable Diffusion has significantly improved the perceptual quality of All-in-One image Restoration (AiOR) methods, while also enhancing their generalization capabilities. However, these LDM-based frameworks suffer from slow inference due to their iterative denoising process, rendering them impractical for time-sensitive applications. Visual autoregressive modeling (VAR), a recently introduced approach for image generation, performs scale-space autoregression and achieves comparable performance to that of state-of-the-art diffusion transformers with drastically reduced computational costs. Moreover, our analysis reveals that coarse scales in VAR primarily capture degradations while finer scales encode scene detail, simplifying the restoration process. Motivated by this, we propose RestoreVAR, a novel VAR-based generative approach for AiOR that significantly outperforms LDM-based models in restoration performance while achieving over $\mathbf{10\times}$ faster inference. To optimally exploit the advantages of VAR for AiOR, we propose architectural modifications and improvements, including intricately designed cross-attention mechanisms and a latent-space refinement module, tailored for the AiOR task. Extensive experiments show that RestoreVAR achieves state-of-the-art performance among generative AiOR methods, while also exhibiting strong generalization capabilities.
@@ 1 Introduction
Image restoration is a complex inverse problem that aims to recover clean images from degradations, such as haze, rain, snow, blur, and low-light conditions. Recently, the paradigm of All-in-One image Restoration (AiOR) has emerged, where a single network is trained to handle multiple degradation types. Existing AiOR methods can be broadly categorized into non-generative and generative approaches. Non-generative models such as AirNet (Li et al., 2022), PromptIR (Potlapalli et al., 2024), InstructIR (Conde et al., 2025), AWRaCLe (Rajagopalan & Patel, 2024), and AdaIR (Cui et al., 2024), deterministically map degraded images to their clean counterparts. While these methods offer fast inference and reliable pixel-level restoration performance, they often fail to generalize to diverse degradations encountered in real-world scenarios. To overcome this challenge, recent works have adopted generative models that aim to capture the distribution of clean images and produce more perceptually realistic outputs. Early works (Chen et al., 2022; Kupyn et al., 2018) based on GANs (Goodfellow et al., 2020) attempted this through adversarial learning, but suffered from mode collapse and unstable training. To improve fidelity and training stability, DiffUIR (Zheng et al., 2024) and DA-CLIP (Luo et al., 2023) employed pixel-space diffusion models (Ho et al., 2020). However, their high computational cost makes large-scale pretraining infeasible, limiting their ability to learn strong generative priors. In contrast, recent methods such as Diff-Plugin(Liu et al., 2024), AutoDIR(Jiang et al., 2023), and PixWizard (Lin et al., 2024) leverage latent diffusion models (LDMs), such as Stable Diffusion (Rombach et al., 2022). By operating in a latent space, LDMs significantly reduce computational costs, enabling large-scale pretraining which equips them with strong generative priors of natural images. These priors allow LDM-based AiOR methods to deliver perceptually realistic restoration and improved generalization to real-world degradations.
@@
Despite their advantages, LDM-based AiOR methods have some shortcomings. (1) LDMs require multiple denoising steps during inference, resulting in significantly longer runtimes compared to non-generative models. Their slow inference speeds pose challenges for applications that demand real-time processing, such as video surveillance or autonomous navigation. (2) LDMs rely on variational autoencoders (VAEs) (Kingma, 2013) which are primarily trained for generative diversity, rather than accurate pixel-level reconstruction. Consequently, the restored images obtained from LDM-based AiOR methods exhibit loss of fine structural details, hindering their performance.
@@
Autoregressive models have driven rapid advances in natural language processing through large language models (LLMs) such as GPT-3 (Radford et al., 2019b) and LLaMA (Touvron et al., 2023). These models generate outputs by predicting the next token, conditioned on previously generated tokens. Recently, Visual AutoRegressive (VAR) Modeling (Tian et al., 2024) introduced scale-space autoregression for image generation, performing next-scale prediction in the latent space of a multi-scale vector-quantized VAE (VQVAE). VAR achieves performance comparable to state-of-the-art diffusion models such as DiT-XL/2 (Peebles & Xie, 2023), while operating $45\times$ faster. Despite its success in generative tasks, the application of VAR to low-level vision tasks such as image restoration remains largely unexplored. To the best of our knowledge, only two prior works—VarSR (Qu et al., 2025) and Varformer (Wang & Zhao, 2024)—have used VAR for image restoration. VarSR focused exclusively on the super-resolution task, while Varformer utilized intermediate VAR features to guide a separate non-generative network for AiOR. In contrast, our approach is generative and fully exploits the strong priors of the pretrained VAR model by training it directly for the AiOR task. Our analysis in Sec. 3.2 also reveals that the scale-space decomposition of VAR captures degradations predominantly in coarse scales and scene-level details in fine scales, making it well-suited for AiOR.
@@
Figure 1: RestoreVAR, our proposed VAR-based (Tian et al., 2024) scale-space generative AiOR model (a), significantly outperforms LDM-based methods as shown in (b). RestoreVAR also offers drastic reductions in computational complexity as shown in (c).
To this end, we introduce RestoreVAR, a novel generative approach for AiOR that addresses some of the key limitations of LDM-based approaches. Firstly, RestoreVAR adopts the autoregressive structure of VAR, achieving state-of-the-art generative AiOR performance with over $\mathbf{10\times}$ faster inference than LDM-based methods (see Fig. 1). Secondly, RestoreVAR employs cross-attention mechanisms conditioned on the degraded image latents, enabling the model to maintain spatial consistency and minimize hallucinations. Thirdly, to mitigate the loss of fine details by the vector quantization and VAE decoding processes, we propose a lightweight (only $\sim\mathbf{3}$% overhead) non-generative latent refinement transformer which predicts de-quantized latents from the outputs of VAR. Additionally, we fine-tune the VAE decoder to operate on these continuous latents, further enhancing reconstruction quality. Finally, through extensive experiments, we demonstrate that RestoreVAR achieves state-of-the-art performance among generative restoration models, while also exhibiting strong generalization to real-world degradations. To summarize, our key contributions are:
@@
    We propose RestoreVAR, the first VAR-based generative AiOR framework that achieves superior performance and a $\mathbf{10\times}$ faster inference than LDM-based methods.
@@ 2.1 Image restoration
Early restoration models primarily addressed specific degradations  (He et al., 2009; Zhang et al., 2020; Wang et al., 2019; Yasarla & Patel, 2019; Zhang et al., 2021a; Nah et al., 2017). Other methods such as Restormer (Zamir et al., 2022), MPRNet (Zamir et al., 2021) and SwinIR (Liang et al., 2021) introduced architectures for any single restoration task. However, they are restricted to handle one degradation at a time, making them ineffective for multiple degradations. All-in-One image Restoration (AiOR) methods aim to tackle multiple corruptions with a single model. Early approaches include non-generative models such as All-in-one (Li et al., 2020) and Transweather (Valanarasu et al., 2022). PromptIR (Potlapalli et al., 2024) used learnable prompts while AWRaCLe (Rajagopalan & Patel, 2024) utilized visual in-context learning to extract degradation characteristics. Other approaches such as InstructIR (Conde et al., 2025) adopted textual guidance, and DCPT (Hu et al., 2025) proposed a novel pre-training strategy for AiOR. DFPIR (Tian et al., 2025) proposed a feature perturbation strategy for AiOR. Recent AiOR methods have adopted diffusion models. Pixel-space diffusion models (PSDMs) such as DA-CLIP (Luo et al., 2023) and DiffUIR (Zheng et al., 2024) demonstrated improved AiOR performance but lacked robust generative priors. Recent methods have utilized the strong priors of LDMs for AiOR. Diff-Plugin (Liu et al., 2024) adopts task plugins to guide an LDM for AiOR. AutoDIR (Jiang et al., 2023) automatically detects and restores degradations using an LDM. PixWizard (Lin et al., 2024) is a multi-task SD-XL (Podell et al., 2023) based model capable of performing AiOR among other tasks. However, LDM-based approaches are slow at inference time—a limitation we aim to overcome using visual autoregressive modeling (VAR).
@@ 2.2 Autoregressive models in vision
Recent works (Van Den Oord et al., 2016; Tian et al., 2024) have extended autoregressive (AR) models to vision and can be categorized as pixel‐space AR (Van Den Oord et al., 2016; Van den Oord et al., 2016; Chen et al., 2018), token‐based AR (Van Den Oord et al., 2017; Yu et al., 2023; Ramesh et al., 2021) and scale‐space AR (Tian et al., 2024; Ren et al., 2024; Guo et al., 2025). Pixel-space AR predicts raw pixels one by one in raster order, as in PixelRNN (Van Den Oord et al., 2016) and PixelCNN++ (Salimans et al., 2017), but is very slow at high resolutions. Token-based AR compresses images into discrete latent codes via vector quantization (e.g., VQ-VAE (Van Den Oord et al., 2017), VQGAN (Esser et al., 2021)) and then models code sequences with transformers (e.g. ImageGPT (Chen et al., 2020)). This trades-off codebook size and transformer capacity against tractability for high-resolution generation. Scale-space AR, as introduced in VAR (Tian et al., 2024), generates latents from coarse to fine scales and matches the quality of Diffusion Transformers (Peebles & Xie, 2023) at a fraction of the inference cost. HART (Tang et al., 2024) scales VAR to higher resolution and uses a MLP-based diffusion refiner to convert discrete VAR latents into continuous representations. Despite VAR’s success in generative tasks, it remains underexplored for image restoration with only two prior works-VarSR (Qu et al., 2025) and Varformer (Wang & Zhao, 2024). VarSR addressed super-resolution, while Varformer used VAR’s features to guide a non-generative AiOR model. In contrast, RestoreVAR is a generative model which directly trains VAR for AiOR.
@@ 3 Proposed Method
We first explain the working principles behind VAR for image generation. We then describe our scale-space analysis of VAR and detail RestoreVAR, our proposed VAR-based approach for AiOR.
@@ 3.1 Preliminaries: Visual Autoregressive Modelling
Visual Autoregressive Modelling, or VAR, is a novel autoregressive class-conditioned image generation method which uses a GPT-2 (Radford et al., 2019a) style decoder-only transformer architecture for next-scale prediction. The VAR transformer operates in the latent space of a multi-scale VQVAE which uses $K$ scales. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, the VQVAE encoder outputs a latent representation $f_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. Hereafter, we will refer to $f_{\text{cont}}$ as the continuous latent, and the latent obtained after quantization as discrete latent. Instead of directly quantizing $f_{\text{cont}}$, a multi-scale residual quantization using a shared codebook across $K$ spatial scales is performed. First, the residual and accumulated quantized (or discrete) reconstruction of $f_{\text{cont}}$ are initialized as $f_{\text{res}}^{(0)}:=f_{\text{cont}}$ and $f_{\text{quant}}^{(0)}:=0$, respectively. At each scale $k=1,\ldots,K$, an index map $r_{k}\in\mathbb{Z}^{H_{k}\times W_{k}}$ is obtained by quantizing the downsampled residual feature:
$r_{k}:=\text{quantize}\left(\text{downsample}\left(f_{\text{res}}^{(k-1)}\right)\right).\vskip-3.0pt$
@@
$h_{k}:=\phi_{k}\left(\text{upsample}\left(e(r_{k})\right)\right),\in\mathbb{R}^{H_{K}\times W_{K}\times C}.\vskip-3.0pt$
@@
$f_{\text{quant}}^{(k)}:=f_{\text{quant}}^{(k-1)}+h_{k},\quad f_{\text{res}}^{(k)}:=f_{\text{cont}}-f_{\text{quant}}^{(k)}.\vskip-3.0pt$
@@
$p(r_{1},r_{2},\ldots,r_{K})=\prod_{k=1}^{K}p(r_{k}\mid r_{1},r_{2},\ldots,r_{k-1}).\vskip-4.0pt\vskip-3.0pt$
@@
3.2 Scale-Space Analysis of VAR
In addition to VAR’s competitive performance to LDMs with far superior inference speed, we found that its residual scale-space decomposition focuses on degradations and scene-level details across different scales. To demonstrate this, we consider clean (GT)–degraded image pairs and compute their scale-wise residual indices $\{r^{\text{GT}}_{k}\}_{k=1}^{K}$ and $\{r^{\text{Deg}}_{k}\}_{k=1}^{K}$, respectively, where $K=10$. We define coarse scales as $k=1,\ldots,5$ (low-resolution index maps) and fine scales as $k=6,\ldots,10$ (higher-resolution index maps). The first and last columns of Fig. 2 show reconstructions from $r^{\text{Deg}}_{k}$ and $r^{\text{GT}}_{k}$, respectively. In column 2, we replace the coarse scales of $r^{\text{GT}}_{k}$ with those from $r^{\text{Deg}}_{k}$. This introduces the degradation, although fine scales remain unchanged. Next, in column 3 we replace the fine scales of $r^{\text{GT}}_{k}$ with those from $r^{\text{Deg}}_{k}$, which yields a clean image with some loss of fine details. These observations indicate that coarse scales in VAR capture degradations, while finer scales encode scene-level detail. Notably, this observation holds across multiple degradations and simplifies restoration for VAR as removing degradations requires correctly predicting only the early scales which contain a small number of tokens, while scene details can be reconstructed in subsequent scales.
@@ 3.3 RestoreVAR
Degraded GT++coarse GT++fine GT
Haze
Refer to caption Refer to caption Refer to caption Refer to caption
Snow
Refer to caption Refer to caption Refer to caption Refer to caption
Rain
Refer to caption Refer to caption Refer to caption Refer to caption
Low-light
Refer to caption Refer to caption Refer to caption Refer to caption
Blur
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: VAR captures degradations in early scales (coarse) and scene-level details in later scales (fine). Degraded and GT are VQVAE reconstructions of the degraded and ground truth images. GT++coarse replaces early GT scales with degraded ones, while GT++fine replaces the late GT scales.
We now describe RestoreVAR, our proposed approach that effectively adapts VAR for AiOR, leveraging its substantial inference speed advantage over LDMs. Given a degraded image $I_{\text{deg}}\in\mathbb{R}^{H\times W\times 3}$, the goal is to predict a clean output $I_{\text{clean}}$, close to the ground-truth $I_{\text{gt}}$. Adapting VAR to AiOR is non-trivial due to the need for high-quality pixel-level reconstruction, which is compromised by two factors: (1) VAR’s strong generative priors can cause hallucinations in the restored images without proper conditioning. (2) Vector quantization and VAE decoding introduce artifacts that hinder pixel-level restoration. RestoreVAR addresses these challenges through architectural enhancements, including cross-attention to incorporate semantic guidance from the degraded image, and a novel non-generative transformer that refines discrete latents into their continuous form to preserve fine details in the restored image. We describe these components below.